Arizona LBTI Commissioning - 07 November 2010 UT

This wiki page includes John's notes on the center bent Gregorian focal station commissioning aspects of LBTI Commissioning. This represents only a small fraction of the total activity.

Observers: JHill, PHinz, Vidhya, Vanessa, Bill
Telescope Operators: DHuerta
Instrument Support: JMorris
Telescope Support: JUrban

Summary

We made 12 microns fringes on several stars with the DX adaptive optics loop closed, and an eyeball active diffraction pattern on the SX side. This ranged from very good seeing before the cold front, to very bad seeing after the cold front.

We believe we have understood that the piezo modulation for the pyramid is working poorly above 200 Hz. Therefore we are achieving poor tip-tilt rejection.

Lost ~2 hours to weather.

Note that LBTI has received honors beyond imagination for last month's fringes. They are now a crossword puzzle word. See http://wordsdomination.com/lbti.html

Details

Adjusted SX M3 Tip=-540 Tilt=-540 trying to improve the LBTI pupil. (Curiously, it has an offset of the M2 image similar to LUCI1.)

Intial Star at EL=72 SX IE=-30 CA=-28 with SX TIp=+19 Tilt=+27
SX IE=-13 CA=-2 with SX Tip=0 Tilt=0

DX IE=+3 CA=-13

All the Tip/Tilts are applied with M1M2 Mode 1 rotation.

Tonight we only required modest (<1000 nm) coma correction compared to the collimations tables.

02:20 IE=+3 CA=-14 SX Tip=-16 Tilt=+11 for both stars.

02:45 Fringes at 12 microns with DX locked and SX showing diffraction rings with eyeball collimation.

04:20 Close for thickening clouds.

05:56 Reopen with scattered thin clouds. T=5.5 degC.

06:10 Found star on SX with IE=-4 CA=-7 plus SX Tip=-16 Tilt=+8

06:25 Found star on DX with IE=-53 CA=-4. Why is this position 56 arcsec away from where we were before in EL?

The evidence seems to indicate that the piezos that modulate the DX star on the pyramid are not working correctly above ~200 Hz. This (combined with various telescope vibrations) apparently gives us poor tip-tilt correction.

08:28 SX at IE=-19 CA=0 with Tip=-16 Tilt=+8

08:32 DX on MIRAC at IE=-58 CA=-1 at EL=68
Translated M1,M2 by +1 X and -1 Y
SX also on MIRAC with SX Tip=+24 Tilt=+14

09:20 IE=-47 CA=0 at EL=62 with SX Tip=+24.5 Tilt=+12

09:40 IE=-12 CA=-13 at EL=74 with SX Tip=-18 Tilt=+25 (The mysterious 56 arcsec have gone away?)

10:30 SX at IE=-21 CA=+12
DX at IE=-50 CA=+4 at EL=44
Both with SX Tip=+29 Tilt=+9

Now that we have 4 data points, perhaps we have a chance at a co-pointing correction to the collimation models. Although it may be that the time order shows more of a trend than the elevation order. Temperature is now dropping rapidly.

Assume IE=0 CA=0 for the pointing model.

EL SX Tip SX Tilt DX Tip DX Tilt
44 -21 +13 -50 +4
62 -22 +12 -47 0
68 -34 +13 -58 -1
74 -30 +12 -12 -13
80 -26 +13 -50 +2

11:02 DX IE=-50 CA=+2 ay EL=80
with SX Tip=+24 Tilt=+11

12:28 Close for industrially bad seeing.

-- JohnHill - 07 Nov 2010
Topic revision: r4 - 07 Nov 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback